Please help me with this equation: Solve 0.02x + 0.5 = 2.6
Answer by sejackson91(8)

0.02x + 0.5 = 2.6
first you subtract 0.5 from both sides
0.02x + 0.5 - 0.5 = 2.6 - 0.5
0.02x = 2.1
then you divide both sides by 0.02
0.02x/0.02 = 2.1/0.02
x = 105